OpenStudy (haseeb96):

A jet fighter is going in 700 km/h .how much distance will it cover in 15 mins ? please tell me how can i solve it ???

OpenStudy (anonymous):

Use this trick: 15mins=1/4 of an hour. You can multiply 700km/h by 1/4h to determine the distance covered in km.

OpenStudy (anonymous):

you can use the kinematic equation X=Xo + Vot + 1/2at^2 X= final position Xo = initial position Vo = initial velocity t= time a=acceleration Since the problem doesn’t mention the jet speeding up or slowing down, you can assume there is no acceleration so a=0 therefore you’re left with X=Xo + Vot you can also set Xo = 0 because you’re trying to find how far this jet traveled from a starting point. So you can set that starting point at 0 km then find out how many km it traveled AFTER traveling 700km/hr for 15 minutes.so now that a=0 and Xo=0 you’re left with X=Vot be sure to convert your velocity (700km/hr) to km/min. after you do that your units for time will be the same and you can find out how far it traveled… or the jet’s “final” position after 15 minutes in this case… Hope that helps.

OpenStudy (haseeb96):

Speed = distance /time 700= distance/ 15 min 700 x 1/4 = distance 175 km is the correct answer ???

OpenStudy (anonymous):

Saeeddiscover's method is way easier for this particular problem. It would be 700km/hr * .25hr (1/4 hr). This way the units for time cancel and you're left with km.